### v1

@@ 8 Conclusion
The inscriptions boom between December 2023 and January 2024 was the first real-life stress test for various EVM-compatible blockchains. Our study of Ethereum, Arbitrum, Base, Optimism, and zkSync Era during this period revealed stark differences between L1 and L2 blockchains and among rollups.
@@
First, we observed that, on specific peak days, transactions related to inscriptions accounted for more than 89% on Arbitrum, more than 88% on zkSync Era, over 53% on Ethereum, and almost 35% on Optimism and Base. The surge in on-chain inscriptions overwhelmed some rollups, causing them to be unprepared for the increased transaction volume. This resulted in downtimes and delays in transaction processing.
@@
We further found that 99% of these inscription transactions across all examined blockchains were driven by the claim or minting of inscription-based meme-coins, followed by a modest amount of trading activity, mostly associated with listing operations on marketplaces. All chains show a dominant token that encompasses more than 70% of all inscription transactions. Interestingly, token variability is greater in Ethereum.
Even though trading of these inscription tokens is limited, we still find some interesting activity. When we look at trading values of the two top tokens on zkSync Era, we see that they range between \$0.1 and  \$0.8. These trades typically occur in batches of 4 or 5 tokens, thus offering minimal compensation to the seller to cover the gas fees required for minting and listing the inscriptions. This explains in part the interest of inscription traders in moving to cheaper chains as the feasibility of trading inscriptions at such low prices is due to the low gas fees on rollups.
@@
Looking at the impact of encryption on gas fees, we concluded that gas fees on rollups either remained stable or decreased during the transaction volume peak. Particularly, the median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and zkSync Era. On the other hand, Ethereum experienced a rise in fees, which is not surprising. Unlike L1 blockchains, rollups are designed to offer lower gas fees as transaction counts increase, due to their ability to compress multiple L2 transactions into a single batch on the L1 network
zkSync Era was particularly efficient in reducing gas fees during the inscriptions boom. ZK rollups, like zkSync Era, spread the cost of generating validity proofs across many transactions. In addition, they only need to post state changes and commitments in the L1 blockchain instead of the full transaction data, further allowing cost sharing in times of high transaction volume. for This results in lower costs per transaction with higher volumes, as observed on specific dates for zkSync Era.
Inscriptions, initially proposed as a workaround for the lack of virtual machines on the Bitcoin blockchain, proved their applicability in EVM blockchains. Nevertheless, the introduction of temporary data storage on Ethereum — blobs — and rollups’ dependence on them — puts the security and longevity of blob- and rollup-based inscription into question. In fact, following an 18-day interval, BlobScriptions are no longer available on the Ethereum blockchain and are solely maintained in the creator protocol’s indexer, which is external to the blockchain. Likewise, rollups are not obligated to retain (beyond blobs) the transaction input data utilized by rollup-based inscriptions, which brings up concerns regarding the ownership of these inscriptions in the long-run.
@@
The future of inscriptions as a stable market for token trading in EVM-chains is still not cemented. The questions about data availability, the still limited trading volume, and the UI challenges of interacting with inscriptions highlight how nascent the market still is. It is nevertheless an interesting trend and a lot of innovation can still occur to turn the space into a more mature one.

### v2

@@ 8 Conclusion
The inscriptions boom between December 2023 and January 2024 was the first real-life stress test for various EVM-compatible blockchains. Our study of Ethereum, Arbitrum, Base, Optimism, and ZKsync Era during this period revealed stark differences between L1, L2blockchains, and among rollups.
@@
First, we observed that, on specific peak days, transactions related to inscriptions accounted for nearly 90% on Arbitrum and ZKsync Era, and over 53% on Ethereum, and almost 35% on Optimism and Base. The surge in on-chain inscriptions overwhelmed some rollups, causing them instability due to the increased transaction volume. This resulted in downtimes and delays in transaction processing.
@@
We further found that 99% of these inscription transactions across all examined blockchains were driven by the claim or minting of inscription-based meme-coins, followed by a modest amount of trading activity, mostly associated with listing operations on marketplaces. All chains show a dominant token that represents more than 74% of all inscription transactions.
@@
Looking at the impact of inscriptions on gas fees, we concluded that gas fees on rollups either remained stable or decreased during the transaction volume peak. Particularly, the median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and ZKsync Era. On the other hand, Ethereum experienced a rise in fees, which is not surprising. Unlike L1blockchains, rollups are designed to offer lower gas fees as transaction counts increase, due to their ability to compress multiple L2transactions into a single batch on the L1network.
@@
Moreover, the future of inscriptions as a stable market for token trading in EVM-chains is still not cemented. For instance, Data Availability (DA)and User Interface (UI)issues can complicate interactions with inscriptions, highlighting how nascent the market still is. It is nevertheless an interesting trend and a lot of innovation can still occur to turn the space into a more mature one.
Finally, to promote scientific reproducibility and facilitate further research, we plan to make our datasets and scripts publicly available.
Acknowledgments
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
Appendix 0.A Glossary
Following is a list of important notations used in this paper.
\@glotype@acronym@title